# GTPO and GRPO-S: Token and Sequence-Level Reward Shaping with Policy Entropy

### v5

@@ 5 Conclusion
In this paper, we addressed the fundamental challenge of coarse-grained credit assignment, a critical flaw in aligning large language models for complex reasoning. We proposed a novel framework centered on dynamic entropy weighting, which introduces two new algorithms: Group Token Policy Optimization (GTPO) for precise, token-level supervision, and a computationally efficient variant, Sequence-Level GRPO (GRPO-S). Our approach repurposes policy entropy as a proxy for model uncertainty to concentrate the learning signal at critical decision points, thereby enabling principled, fine-grained credit assignment. Extensive experiments demonstrate that our methods consistently outperform strong DAPO and GRPO baselines across multiple reasoning benchmarks, confirming the efficacy of the proposed entropy-weighting mechanism. Ultimately, our findings suggest that harnessing and directing model uncertainty is a promising frontier for developing the next generation of powerful and reliable AI systems, as demonstrated qualitatively in the case study in Appendix E.
@@ A.1 The Evolution of LLM Alignment Algorithms
LLM alignment techniques aim to make model behavior conform to human expectations and values. The field was initially dominated by PPO-based RLHF. This classic paradigm consists of three stages: supervised fine-tuning (SFT), reward model training, and reinforcement learning optimization. Despite its power, its process is complex, sensitive to hyperparameters, and often unstable during training. To overcome these challenges, the research community has shifted towards more direct optimization methods. Direct Preference Optimization (DPO) was a landmark work that cleverly transformed the reward maximization problem into a simple classification loss, completely bypassing explicit reward modeling and the RL process [Rafailov et al., 2023]. The success of DPO has spawned a series of variants, such as ODPO [Amini et al., 2024], which considers the strength of preferences, and Preference Tuning LLMs with TRL [Hugging Face, 2024], which aims to solve overfitting, collectively advancing the RL-free alignment paradigm.
@@ A.2 The Rise of Value-Function-Free Policy Optimization
Our work builds directly on value-function-free policy optimization methods. Group Relative Policy Optimization (GRPO), introduced by DeepSeekMath, is a representative of this direction [Shao et al., 2024]. The core mechanism of GRPO is: for a given prompt, sample a group of $G$ sequences from the current policy, and then use the average reward within this group as a baseline to calculate the advantage for each sequence [Kilcher, 2024]. This design eliminates the need for a separate value function network, greatly reducing memory consumption and computational complexity, which has led to great success in tasks like mathematical reasoning. However, the original GRPO is also sensitive to reward noise and can be unstable during training, which has prompted subsequent research for improvements.
@@ A.3 A Technical Comparison with the DAPO Baseline
Decoupled Clip and Dynamic sAmpling Policy Optimization (DAPO) [Yu et al., 2025] is currently the state-of-the-art (SOTA) method for GRPO-style training in the open-source community. DAPO significantly improves the performance and stability of GRPO by introducing four key techniques: 1) Clip-Higher: Encourages model exploration and prevents entropy collapse by relaxing the upper bound of the PPO clipping range. 2) Dynamic Sampling: Filters out sample groups that are either all successful or all failures, ensuring that each training batch contains effective gradient signals, thus improving training efficiency. 3) Token-Level Policy Gradient Loss: A core improvement of DAPO, its objective function averages the loss over all tokens in a batch, rather than first summing within a sequence and then averaging across sequences as in the original GRPO [Yu et al., 2025]. 4) Overlong Reward Penalty: Penalizes excessively long generated sequences to reduce reward noise.
@@ A.4 Entropy as a Heuristic for Cognitive Effort in LLMs
Using model entropy as a measure of uncertainty has a long history in the machine learning field. Recent research has shown that during the reasoning process of LLMs, the entropy of the model’s generated probability distribution is highly correlated with cognitive uncertainty. For example, Cheng et al. [2025] found that in successful reasoning paths, high-entropy regions often correspond to steps where the model engages in meaningful exploration and critical logical reasoning. This finding provides strong support for our use of entropy as a heuristic for credit assignment and forms the cornerstone of our methodology.
@@ B.2 Unifying the objective function of GRPO at Token-Level
Next, the $\text{mean}(\{R_{i}\}_{i=1}^{G})$ part of the GRPO objective function is analyzed. First, the ideal state of the advantage function is known to be A=QVA=Q-V, where $Q$ is the action-value and $V$ is the state-value. Therefore, during sampling, Q and V need to be estimated as accurately as possible. The analysis in A.1 for $Y$ is actually an analysis of the estimation method for $Q$. DAPO has already modified this, but changing the estimation method for $V$ can be considered. The term $\text{mean}(\{R_{i}\}_{i=1}^{G})$ is the estimate for $V$. Theoretically, there is a more accurate estimation method, which is proven below.
@@
B.3 Analysis of the Sequence-Level Objective (GRPO-S)
A parallel analysis applies to the sequence-level objective function. Since the reward for incorrect sequences only affects the gradient’s magnitude, not its direction, the analysis similarly centers on the reward redistribution for correct sequences. The reshaping of sequence-level rewards, $\hat{r}_{i}^{+}$, is designed to preserve the total reward across the batch. This is confirmed by the following relationship:
$\sum_{i=1}^{G}\hat{r}_{i}^{+}=(\beta_{1}+\beta_{2})\sum_{i=1}^{G}r_{i}\approx\sum_{i=1}^{G}r_{i},$
where $\beta_{1}+\beta_{2}\approx 1$ is set. From this property, it follows that the expected mean reward is conserved:
$\mathbb{E}[\text{mean}(\{\hat{r}_{k}^{+}\}_{k=1}^{G})]\approx\mathbb{E}[\text{mean}(\{r_{k}\}_{k=1}^{G})].$
This conservation property ensures that the expectation of our modified advantage function, $\mathbb{E}[\hat{A}_{i}^{+}]$, remains approximately equal to the expectation of the advantage function from the original GRPO algorithm, $\mathbb{E}[\hat{A}_{i}]$ ($\mathbb{E}[\hat{A}_{i}^{+}]\approx\mathbb{E}[\hat{A}_{i}]$). Consequently, with the entropy term detached from the gradient, the expected policy gradient of our sequence-level objective, $\mathcal{J}_{\text{GRPO-S}}(\theta)$, is approximately equivalent to that of the baseline GRPO objective:
$\mathbb{E}[\nabla_{\theta}\mathcal{J}_{\text{GRPO-S}}(\theta)]\approx\mathbb{E}[\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)].$
Following a similar derivation to the token-level case, it is concluded that our sequence-level reward shaping preserves the convergence properties of the underlying reinforcement learning algorithm. The modification primarily serves to reshape the reward landscape to influence training dynamics, such as by reducing gradient variance, without altering the fundamental optimization direction in expectation.
B.4 Gradient
This section briefly analyzes the expected gradient of the GTPO and GRPO-S objective function. Since the entropy term is detached during gradient calculation, it is easy to calculate that
$\begin{split}\nabla_{\theta}J_{\text{GTPO}}(\theta)=\mathbb{E}\bigg[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}w_{i,t}(\theta)\tilde{A}_{i,t}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})\bigg].\end{split}$ (16)
And according to policy gradient theorem, it is easy to calculate that
$\begin{split}\nabla_{\theta}J_{\text{GRPO-S}}(\theta)=\mathbb{E}\bigg[\frac{1}{G}\sum_{i=1}^{G}\widehat{w}_{i}(\theta)\widehat{A}_{i}\cdot\nabla_{\theta}\log\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})\cdot\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\log\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})\bigg].\end{split}$ (17)

### v6

@@ 3.3 Reward Trajectories and Sample Efficiency
Figure 4: Comparison of three different methods.
Figure 5: Hyperparameter comparison.
The mean reward trajectories on the test set (Fig. 3) illuminate the benefits of our approach, demonstrating that our methods not only achieve a significantly higher final reward ceiling but do so with strong sample efficiency. The models largely converge within 210 training steps, a finding substantiated by the training set reward curves, which are deferred to Appendix E.4 for brevity. This rapid convergence indicates that the benefits of enhanced exploration do not come at the cost of slower learning.
Our methods foster sustained exploration and prevent policy collapse, establishing a clear causal chain from our entropy-weighted reward to superior performance. A detailed analysis and the empirical evidence for this mechanism are presented in Appendix E.1.
3.4 Hyperparameter Sensitivity Analysis
Reward shaping adds extra degrees of freedom that control the trade-off between exploration and optimizing the primary objective. To test this balance between exploration and exploitation. We conduct a sensitivity analysis on the key reward shaping hyperparameters for GTPO and GRPO-S, with results presented in Fig. 5. Across all tested configurations, both of our methods demonstrate robust and significant performance gains over the DAPO baseline on Mean@32 and Pass@32 metrics. Among them, GRPO-S exhibits higher stability across settings, providing a clear view of the performance dynamics. As the weight of the entropy bonus ($\beta_{2}$) increases from 0.1 to 0.2, we observe a clear decline in overall performance for GRPO-S. This finding confirms that while exploration is beneficial, an excessive entropy bonus can detract from optimizing the primary task objective.
4 Discussion
Future Work. The concept of relative entropy itself warrants deeper exploration. While this work highlights its importance, the optimal method for comparison remains an open question. Three distinct approaches for measuring this relativity are identified. Visualizing the token entropies as a matrix $H$, where $H_{i,t}$ is the entropy of the $t$-th token in the $i$-th response, these methods can be understood as: a column-wise comparison, which compares the entropy of tokens at the same position across different responses (i.e., within a column of $H$); a row-wise comparison, which compares tokens at different positions within the same response (i.e., within a row of $H$); and a matrix-wise comparison, which compares all tokens from all responses collectively (i.e., across the entire matrix $H$). Further experimentation is needed to verify which of these approaches is superior. The conceptual difference is illustrated in Fig. 4.
@@ 5 Conclusion
In this paper, we address the fundamental challenge of coarse-grained credit assignment, a critical flaw in aligning large language models for complex reasoning. We propose a novel framework centered on dynamic entropy weighting, which introduces two new algorithms: Group Token Policy Optimization (GTPO) for precise, token-level supervision, and a computationally efficient variant, Sequence-Level GRPO (GRPO-S). Our approach repurposes policy entropy as a proxy for model uncertainty to concentrate the learning signal at critical decision points, thereby enabling principled, fine-grained credit assignment. Extensive experiments demonstrate that our methods consistently outperform strong DAPO and GRPO baselines across multiple reasoning benchmarks, confirming the efficacy of the proposed entropy-weighting mechanism. Ultimately, our findings suggest that harnessing and directing model uncertainty is a promising frontier for developing the next generation of powerful and reliable AI systems, as demonstrated qualitatively in the case study in Appendix F.
@@ A.1 The Evolution of LLM Alignment Algorithms
LLM alignment techniques aim to make model behavior conform to human expectations and values. The field is initially dominated by PPO-based RLHF. This classic paradigm consists of three stages: supervised fine-tuning (SFT), reward model training, and reinforcement learning optimization. Despite its power, its process is complex, sensitive to hyperparameters, and often unstable during training. To overcome these challenges, the research community shifts towards more direct optimization methods. Direct Preference Optimization (DPO) is a landmark work that cleverly transformed the reward maximization problem into a simple classification loss, completely bypassing explicit reward modeling and the RL process (Rafailov et al., 2023). The success of DPO spawns a series of variants, such as ODPO (Amini et al., 2024), which considers the strength of preferences, and Preference Tuning LLMs with TRL (Hugging Face, 2024), which aims to solve overfitting, collectively advancing the RL-free alignment paradigm.
@@ A.2 The Rise of Value-Function-Free Policy Optimization
Our work builds directly on value-function-free policy optimization methods. Group Relative Policy Optimization (GRPO), introduced by DeepSeekMath, is a representative of this direction (Shao et al., 2024). The core mechanism of GRPO is: for a given prompt, sample a group of $G$ sequences from the current policy, and then use the average reward within this group as a baseline to calculate the advantage for each sequence (Kilcher, 2024). This design eliminates the need for a separate value function network, greatly reducing memory consumption and computational complexity, which leds to great success in tasks like mathematical reasoning. However, the original GRPO is also sensitive to reward noise and can be unstable during training, which prompts subsequent research for improvements.
@@ A.3 A Technical Comparison with the DAPO Baseline
Decoupled Clip and Dynamic sAmpling Policy Optimization (DAPO) (Yu et al., 2025) is currently the state-of-the-art (SOTA) method for GRPO-style training in the open-source community. DAPO significantly improves the performance and stability of GRPO by introducing four key techniques: 1) Clip-Higher: Encourages model exploration and prevents entropy collapse by relaxing the upper bound of the PPO clipping range. 2) Dynamic Sampling: Filters out sample groups that are either all successful or all failures, ensuring that each training batch contains effective gradient signals, thus improving training efficiency. 3) Token-Level Policy Gradient Loss: A core improvement of DAPO, its objective function averages the loss over all tokens in a batch, rather than first summing within a sequence and then averaging across sequences as in the original GRPO (Yu et al., 2025). 4) Overlong Reward Penalty: Penalizes excessively long generated sequences to reduce reward noise.
@@ A.4 Entropy as a Heuristic for Cognitive Effort in LLMs
Using model entropy as a measure of uncertainty has a long history in the machine learning field (Wang et al., 2025c). Recent research shows that during the reasoning process of LLMs, the entropy of the model’s generated probability distribution is highly correlated with cognitive uncertainty. For example, Cheng et al. (2025) find that in successful reasoning paths, high-entropy regions often correspond to steps where the model engages in meaningful exploration and critical logical reasoning. This finding provides strong support for our use of entropy as a heuristic for credit assignment and forms the cornerstone of our methodology.
@@ Appendix B Theoretical Analysis and Proofs
For the purpose of theoretical variance analysis, we assume the reward signals associated with individual tokens are independent and identically distributed (i.i.d.). While token generation is auto-regressive, this simplification allows us to formally derive the variance reduction properties of the global mean estimator.
@@ B.2 Unifying the objective function of GRPO at Token-Level
Next, the $\text{mean}(\{R_{i}\}_{i=1}^{G})$ part of the GRPO objective function is analyzed. First, the ideal state of the advantage function is known to be A=QVA=Q-V, where $Q$ is the action-value and $V$ is the state-value. Therefore, during sampling, Q and V need to be estimated as accurately as possible. The analysis in A.1 for $Y$ is actually an analysis of the estimation method for $Q$. DAPO already modify this, but changing the estimation method for $V$ can be considered. The term $\text{mean}(\{R_{i}\}_{i=1}^{G})$ is the estimate for $V$. Theoretically, there is a more accurate estimation method, which is proven below.
@@
B.3 Theoretical Analysis of Reward Conservation and Convergence
In this section, we provide a formal proof demonstrating that the Dynamic Entropy Weighting mechanism in GTPO preserves the total positive reward mass of the original GRPO objective. This conservation property ensures that the expected policy gradient direction remains consistent with the ground-truth reward maximization, thereby guaranteeing that the positive reward component of GTPO optimizes towards the same global objective as the baseline.
Proposition B.3 (Conservation of Positive Reward Mass).
[…]
$\sum_{i\in O^{+}_{t}}\tilde{r}^{+}_{i,t}=\sum_{i\in O^{+}_{t}}r_{i}.$ (22)
Proof.
Recall the definition of the shaped token reward for a successful sequence $o_{i}\in O^{+}$ at timestep $t$, as provided in the GTPO formulation:
$\tilde{r}^{+}_{i,t}=\alpha_{1}r_{i}+\alpha_{2}\frac{H_{i,t}}{\sum_{k\in O^{+}_{t}}H_{k,t}}\cdot d_{t},$ (23)
where $r_{i}=1$ for all successful sequences, $H_{i,t}$ is the policy entropy, and the summation in the denominator is taken over all active successful sequences at timestep $t$.
We compute the sum of shaped rewards over all sequences $i\in O^{+}_{t}$:
$\sum_{i\in O^{+}_{t}}\tilde{r}^{+}_{i,t}$ $=\sum_{i\in O^{+}_{t}}\left(\alpha_{1}\cdot 1+\alpha_{2}\frac{H_{i,t}}{\sum_{k\in O^{+}_{t}}H_{k,t}}\cdot d_{t}\right)$ (24)
$=\sum_{i\in O^{+}_{t}}\alpha_{1}+\sum_{i\in O^{+}_{t}}\left(\alpha_{2}d_{t}\frac{H_{i,t}}{\sum_{k\in O^{+}_{t}}H_{k,t}}\right).$ (25)
The first term simplifies directly since there are $d_{t}$ terms:
$\sum_{i\in O^{+}_{t}}\alpha_{1}=\alpha_{1}d_{t}.$ (26)
For the second term, we factor out the constants $\alpha_{2}$ and $d_{t}$, observing that the sum of the normalized entropies is unity:
$\sum_{i\in O^{+}_{t}}\left(\alpha_{2}d_{t}\frac{H_{i,t}}{\sum_{k\in O^{+}_{t}}H_{k,t}}\right)=\alpha_{2}d_{t}\frac{\sum_{i\in O^{+}_{t}}H_{i,t}}{\sum_{k\in O^{+}_{t}}H_{k,t}}=\alpha_{2}d_{t}\cdot 1=\alpha_{2}d_{t}.$ (27)
Combining these results:
$\sum_{i\in O^{+}_{t}}\tilde{r}^{+}_{i,t}=\alpha_{1}d_{t}+\alpha_{2}d_{t}=(\alpha_{1}+\alpha_{2})d_{t}.$ (28)
Given the constraint $\alpha_{1}+\alpha_{2}=1$, we arrive at:
$\sum_{i\in O^{+}_{t}}\tilde{r}^{+}_{i,t}=d_{t}=\sum_{i\in O^{+}_{t}}r_{i}.$ (29)
Since the original reward $r_{i}$ is uniformly $1$ for all successful sequences, its sum is exactly the count $d_{t}$. This proves that the total positive reward mass distributed at each timestep is strictly conserved.
Proposition B.4 (Asymptotic Consistency).
Let $\Delta\mathcal{J}(\theta)=\nabla\mathcal{J}_{\text{GTPO}}^{+}(\theta)-\nabla\mathcal{J}_{\text{GRPO}}(\theta)$ be the gradient bias introduced by the entropy-weighted reward shaping, where $\mathcal{J}_{\text{GTPO}}^{+}(\theta)$ is the positive reward component of GTPO. Assume the policy $\pi_{\theta}$ satisfies the Entropy Consolidation Condition: as training iteration $k\to\infty$, the variation in token entropy among successful sequences diminishes, i.e., for any two successful sequences $o_{i},o_{j}\in O^{+}$, $\lim_{k\to\infty}\frac{H_{i,t}}{H_{j,t}}=1$ (almost surely, assuming regularization $\epsilon>0$ prevents singularity). Then, the gradient bias vanishes asymptotically:
$\lim_{k\to\infty}\|\Delta\mathcal{J}(\theta_{k})\|=0.$ (30)
Proof.
Recall that the GTPO reward $\tilde{r}_{i,t}$ is defined as:
$\tilde{r}_{i,t}=\alpha_{1}r_{i}+\alpha_{2}\frac{H_{i,t}}{\sum_{p\in O^{+}_{t}}H_{p,t}}d_{t}.$ (31)
Since $r_{i}=1$ for all successful sequences, the original GRPO reward is constant $1$. The deviation term (exploration bias) for a specific sequence $o_{i}$ is:
$\delta_{i,t}=\tilde{r}_{i,t}-r_{i}=\alpha_{1}+\alpha_{2}d_{t}\frac{H_{i,t}}{\sum_{p=1}^{d_{t}}H_{p,t}}-1.$ (32)
Substituting $\alpha_{1}=1-\alpha_{2}$:
$\delta_{i,t}=(1-\alpha_{2})+\alpha_{2}d_{t}\frac{H_{i,t}}{\sum_{p}H_{p,t}}-1=\alpha_{2}\left(\frac{d_{t}H_{i,t}}{\sum_{p}H_{p,t}}-1\right).$ (33)
Let $\bar{H}_{t}=\frac{1}{d_{t}}\sum_{p\in O^{+}_{t}}H_{p,t}$ be the arithmetic mean of entropies for successful sequences at step $t$. The term can be rewritten as:
$\delta_{i,t}=\alpha_{2}\left(\frac{H_{i,t}}{\bar{H}_{t}}-1\right).$ (34)
Under the Entropy Consolidation assumption, as $k\to\infty$, the entropies of successful paths converge to a uniform level (typically approaching 0 due to policy collapse, or stabilizing at a intrinsic uncertainty level). In either case, the ratio of individual entropy to the group mean approaches unity:
$\lim_{k\to\infty}\frac{H_{i,t}}{\bar{H}_{t}}=1.$ (35)
Consequently, the reward deviation vanishes:
$\lim_{k\to\infty}\delta_{i,t}=\alpha_{2}(1-1)=0.$ (36)
The gradient difference is the expectation of this deviation weighted by the score function:
$\Delta\mathcal{J}(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[\sum_{t}\delta_{i,t}\nabla\log\pi_{\theta}(o_{i,t}|\cdot)\right].$ (37)
Since $\delta_{i,t}\to 0$ uniformly for all successful sequences, by the linearity of expectation (and assuming bounded gradients), we conclude:
$\lim_{k\to\infty}\Delta\mathcal{J}(\theta_{k})=0.$ (38)
This confirms that while the positive reward component of GTPO encourages exploration during the transient phase where entropy variance is high, it asymptotically reduces to the standard GRPO objective, ensuring valid convergence.
B.4 Analysis of GRPO-S
Analogous to the theoretical analysis of GTPO in Appendix B.3, we now present the analysis for GRPO-S. Similar to the token-level analysis, we establish that the sequence-level reward shaping in GRPO-S preserves the fundamental properties of the baseline estimator: reward mass conservation and asymptotic gradient consistency.
Proposition B.5 (Sequence-Level Reward Conservation).
Let $O^{+}=\{o_{1},\dots,o_{n}\}$ be the set of $n$ successful sequences in a sampled group. Let $\hat{H}_{i}$ denote the average entropy of sequence $o_{i}$. Under the GRPO-S reward shaping defined in Eq. (9), if $\beta_{1}+\beta_{2}=1$, then the sum of shaped positive sequence rewards equals the sum of original rewards:
$\sum_{i=1}^{n}\hat{r}^{+}_{i}=\sum_{i=1}^{n}r_{i}=n.$ (39)
Proof.
The GRPO-S shaped reward for a successful sequence $o_{i}$ is given by:
$\hat{r}^{+}_{i}=\beta_{1}r_{i}+\beta_{2}\frac{\hat{H}_{i}}{\sum_{k=1}^{n}\hat{H}_{k}}\cdot n,$ (40)
where $r_{i}=1$ for all $o_{i}\in O^{+}$. Summing over all $n$ successful sequences:
$\sum_{i=1}^{n}\hat{r}^{+}_{i}$ $=\sum_{i=1}^{n}\left(\beta_{1}\cdot 1+\beta_{2}n\frac{\hat{H}_{i}}{\sum_{k=1}^{n}\hat{H}_{k}}\right)$ (41)
$=n\beta_{1}+\beta_{2}n\sum_{i=1}^{n}\frac{\hat{H}_{i}}{\sum_{k=1}^{n}\hat{H}_{k}}.$ (42)
Observing that the summation term $\sum_{i=1}^{n}\frac{\hat{H}_{i}}{\sum_{k=1}^{n}\hat{H}_{k}}=\frac{\sum\hat{H}_{i}}{\sum\hat{H}_{k}}=1$, we have:
$\sum_{i=1}^{n}\hat{r}^{+}_{i}=n\beta_{1}+n\beta_{2}=n(\beta_{1}+\beta_{2}).$ (43)
With the constraint $\beta_{1}+\beta_{2}=1$, it follows that $\sum\hat{r}^{+}_{i}=n$, which matches the sum of the ground-truth rewards $\sum r_{i}$. Thus, the total positive reward signal used for the advantage calculation is conserved at the group level.
Proposition B.6 (Asymptotic Gradient Consistency of GRPO-S).
Let $\mathcal{J}_{\text{GRPO}}(\theta)$ and $\mathcal{J}_{\text{GRPO-S}}(\theta)$ be the expected objective functions under the original and shaped rewards, respectively. Assume the following regularity conditions:
  1. 1.
    Boundedness: The policy score function $\|\nabla_{\theta}\log\pi_{\theta}(o)\|$ is bounded by a constant $M$ for all feasible $o$.
  2. 2.
    Entropy Consolidation: As training progresses ($k\to\infty$), the relative entropy variance among successful sequences vanishes: $\sup_{i\in O^{+}}|\frac{\hat{H}_{i}}{\bar{H}}-1|\xrightarrow{p}0$.
Then, the gradient of the GRPO-S objective converges to the GRPO gradient in $L_{1}$ norm:
$\lim_{k\to\infty}\|\nabla\mathcal{J}_{\text{GRPO-S}}^{+}(\theta_{k})-\nabla\mathcal{J}_{\text{GRPO}}(\theta_{k})\|_{L_{1}}=0.$ (44)
Proof.
Let $O=\{o_{1},\dots,o_{G}\}$ be the sampled group. The gradient estimator for GRPO is:
$g(\theta)=\frac{1}{G}\sum_{i=1}^{G}A_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i}),\quad\text{where }A_{i}=\frac{r_{i}-\mu}{\sigma}.$ (45)
Similarly, the gradient estimator for GRPO-S is:
$\hat{g}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i}),\quad\text{where }\hat{A}_{i}=\frac{\hat{r}_{i}-\hat{\mu}}{\hat{\sigma}}$. (46)
Here, $\mu,\sigma$ are the mean and standard deviation of the original rewards $\{r\}$, and $\hat{\mu},\hat{\sigma}$ are of the shaped rewards $\{\hat{r}\}$.
Step 1: Convergence of the Reward Signal. Recall from Proposition 39 that the shaped reward is $\hat{r}_{i}=\beta_{1}r_{i}+\beta_{2}\frac{\hat{H}_{i}}{\bar{H}}r_{i}$ (since $r_{i}=0$ for incorrect answers, this form holds generally if we define deviation on the correct subset). The deviation is $\delta_{i}=\hat{r}_{i}-r_{i}=\beta_{2}r_{i}(\frac{\hat{H}_{i}}{\bar{H}}-1)$. Under the Entropy Consolidation assumption, for any $\epsilon>0$, there exists a step $K$ such that for all $k>K$, $|\frac{\hat{H}_{i}}{\bar{H}}-1|<\epsilon$. Thus, the shaped rewards converge pointwise to the original rewards:
$\lim_{k\to\infty}|\hat{r}_{i}-r_{i}|=0.$ (47)
Step 2: Continuity of the Advantage Function. The Advantage function $A(r)=\frac{r-\text{mean}(r)}{\text{std}(r)}$ is a continuous mapping with respect to the input vector $\mathbf{r}\in\mathbb{R}^{G}$, provided $\text{std}(r)>0$ (which is true in practice as we sample diverse groups or use $\epsilon$-regularization). Since $\hat{r}_{i}\to r_{i}$ for all $i$, by the Continuous Mapping Theorem:
$\lim_{k\to\infty}|\hat{\mu}-\mu|=0\quad\text{and}\quad\lim_{k\to\infty}|\hat{\sigma}-\sigma|=0.$ (48)
Consequently, the shaped advantage converges to the original advantage:
$\lim_{k\to\infty}|\hat{A}_{i}-A_{i}|=0.$ (49)
Step 3: Convergence of the Expected Gradient. Consider the norm of the difference between the expected gradients:
$\|\Delta\mathcal{J}\|_{L_{1}}$ $:=\|\nabla\mathcal{J}_{\text{GRPO-S}}^{+}(\theta_{k})-\nabla\mathcal{J}_{\text{GRPO}}(\theta_{k})\|_{L_{1}}$ (50)
$=\left\|\mathbb{E}_{O\sim\pi}\left[\frac{1}{G}\sum_{i=1}^{G}(\hat{A}_{i}-A_{i})\nabla_{\theta}\log\pi_{\theta}(o_{i})\right]\right\|$ (51)
$\leq\mathbb{E}_{O\sim\pi}\left[\frac{1}{G}\sum_{i=1}^{G}|\hat{A}_{i}-A_{i}|\cdot\|\nabla_{\theta}\log\pi_{\theta}(o_{i})\|\right].$ (52)
By the Boundedness assumption, $\|\nabla\log\pi\|\leq M$. Thus:
$\|\Delta\mathcal{J}\|_{L_{1}}\leq M\cdot\mathbb{E}_{O\sim\pi}\left[\frac{1}{G}\sum_{i=1}^{G}|\hat{A}_{i}-A_{i}|\right].$ (53)
Let $X_{k}=\frac{1}{G}\sum|\hat{A}_{i}-A_{i}|$ be the random variable representing the average advantage deviation at step $k$. We established in Step 2 that $X_{k}\xrightarrow{a.s.}0$. Since the rewards and advantages are bounded (clipped in implementation), $X_{k}$ is dominated by a bounded constant. By the Lebesgue Dominated Convergence Theorem, we can interchange the limit and the expectation:
$\lim_{k\to\infty}\|\Delta\mathcal{J}\|_{L_{1}}\leq M\cdot\mathbb{E}\left[\lim_{k\to\infty}X_{k}\right]=M\cdot\mathbb{E}[0]=0.$ (54)
This proves that the gradient of the positive reward component of GRPO-S is an asymptotically consistent estimator of the GRPO gradient.
Based on Propositions 39 & 44, and following a proof analogous to that of Theorem 2.4, we derive the following theorem:
Theorem B.7 (The Same Global Optimum - GRPO-S).
GRPO-S shares the same global optimum as DAPO.
Remark B.8.
Note that while Proposition 39 assumes $\beta_{1}+\beta_{2}=1$, this condition is not strictly necessary. The proofs remain valid even if $\beta_{1}+\beta_{2}\neq 1$, as the term $\beta_{1}+\beta_{2}$ merely acts as a constant scaling factor that does not affect the validity of the derivation. In our experimental setup, the sum $\beta_{1}+\beta_{2}$ is configured to be close to 1, but strict equality is not enforced.
Appendix C Alternative Definitions via Geometric Mean
Recall that we used Eq.(11) to define the sequence-level importance weight $\hat{w}_{i}(\theta)$, which is an arithmetic mean. In fact, we can define the sequence-level importance weight via geometric mean: $\hat{w}_{i}(\theta)=(\prod_{t=1}^{|o_{i}|}w_{i,t}(\theta))^{1/|o_{i}|}.$ We can employ the geometric mean as a low-variance proxy for the sequence-level importance weight. Although this introduces bias, it acts as a geometric baseline that stabilizes training for long sequences. Similarly, we have alternative definitions for other key concepts via geometric mean:
$\tilde{r}_{i,t}^{+}$ $=\alpha_{1}r_{i}+\alpha_{2}\frac{H_{i,t}}{(\prod_{k=1}^{n}H_{k,t})^{1/d_{t}}},$ $\tilde{r}_{j,t}^{-}=\alpha_{1}(-1)+\alpha_{2}\frac{1/H_{j,t}}{(\prod_{k=1}^{m}1/H_{k,t})^{1/h_{t}}}(-1),$
$\hat{r}_{i}^{+}$ $=\beta_{1}r_{i}+\beta_{2}\frac{\hat{H}_{i}}{(\prod_{k=1}^{n}\hat{H}_{k})^{1/n}},$ $\hat{r}_{j}^{-}=\beta_{1}(-1)+\beta_{2}\frac{1/\hat{H}_{j}}{(\prod_{k=1}^{m}1/\hat{H}_{k})^{1/m}}(-1),$
Crucially, we note a specific distinction here: if a sequence $o_{*}$ has length less than $t$, its $H_{*,t}$ or $1/H_{*,t}$ is treated as $1$, instead of 0, to ensure the denominator remains non-zero.